Imagine you want to add a sprinkler system and the location of one section of the sprinkler line can be described by the equation
y=-1/2x-4
Complete the table for this equation.
x y (x,y)
___ -1
-2 ___
___ -4
2 ___
8 ___

Answer by stanbon(75887)

Imagine you want to add a sprinkler system and the location of one section of the sprinkler line can be described by the equation
y=-1/2x-4
Complete the table for this equation.
x y (x,y)
___ -1
y = (-1/2)x - 4
---
-1 = (-1/2)x - 4
3 = (-1/2)x
x = -6
--------------------------------
y=(-1/2)x-4
-2 ___
y = (-1/2)(-2) - 4
y = 1-4
y = -3
---------------------------------
y=(-1/2)x-4
___ -4
-4 = (-1/2)x -4
0 = (-1/2)x
x = 0
---------------------------------
y=-1/2x-4
2 ___
y = (-1/2)2 -4
y = -1-4
y = -5
---------------------------------
y=(-1/2)x-4
8 ___
y = (-1/2)8-4
y = -8
====================================
Cheers,
Stan H.